how many different pairs of positive even integers have a product of 120?
Answer by richwmiller(17219)

120=120*1 1 is not even
120=60*2
120=40*3 3 is not even
120=30*4
120=24*5 5 is not even
120-20*6
120-15*8 15 is not even
120=12*10
120=1*3*2*2*2*5
make any two groups of these factors where 1 3 and 5 are not alone nor combined together